# Reinforce Lifelong Interaction Value of User-Author Pairs for Large-Scale Recommendation Systems

### v1

@@ 1. Introduction
2. PROBLEM FORMULATION
Existing RL methods in RS often model user behaviors as infinite request-level markov decision process (MDP) (Sutton, 2018). Specifically, the time interval between adjacent states $\Delta=1$ always holds. However, under the UA interaction space, the time interval between the same UA pair’s adjacent interactions satisfies $\Delta\geq 1$. For a specific UA pair, the interactions are usually sparse due to the large scale of candidate recommended items.
Therefore, we define a novel sparse cross-request interaction MDP (SCRI-MDP) to model the LIV of UA pairs. Formally, it is represented by a tuple of five elements $<\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma>$:
  • State space $\mathcal{S}$: The state $\mathbf{s}_{ua}\in\mathcal{S}$ includes user $u$’s static features (e.g., user ID, location, gender), item and author $a$’s static features (e.g., item ID, author ID) and dynamic features of the $ua$ pair (e.g. cumulative gift count, cumulative watch time).
  • Action space $\mathcal{A}$: The action space $\mathcal{A}$ is defined as the candidate items in the ranking stage, and typically contains a few hundred items. The corresponding action $a\in\mathcal{A}$ is a recommended item list.
  • State transition probability distribution $P$: It is denoted as $P(\mathbf{s}_{ua}^{\prime}|\mathbf{s}_{ua},a)$ determined by the environment, where $\mathbf{s}_{ua}^{\prime}$ represents the next state of the same $ua$ pair.
  • Reward function $\mathcal{R}$: To model the progressive changes of UA relationship, several functions are designed to model the long-term reward of different immediate feedback between $ua$. Let $r_{ua,c},r_{ua,w},r_{ua,f},r_{ua,g}$ be the reward function of click, watch time, follow and gift labels respectively. $C=\{c,w,f,g\}$ denotes the target label set and $n=4$ denotes the cardinality of $C$.
  • Discount factor $\gamma$: $\gamma\in[0,1]$ is used to trade off instant rewards and future rewards.
We introduce a multi-task critic learning (MTCL) architecture to model the LIV of UA pairs by capturing the progressive changes of lifelong UA relationship. There are $n$ critic networks $Q_{\phi_{k}},k=1,...,n$ with $\phi_{k}$ as their trainable parameters to optimize the corresponding cumulative rewards. The final objective function is as follows:
(1) $\max_{\phi_{1},\cdots,\phi_{n}}\sum_{l\in C}\mathbb{E}\left[\sum_{t=0}^{\infty}{r_{ua,\,l}^{t}}\right]$
Each critic network is designed to learn from the states of UA pairs and corresponding rewards. Specifically, at one request of user $u$, RS recommends author $a$’s item $I_{i}$ which is the $t$-th interaction of $ua$, then the LIV of $ua$ pair is defined as:
(2) $Q(\mathbf{s}^{t}_{ua},I_{i})=r^{t}_{ua}+\gamma\max_{j\in\mathcal{I}}Q(\mathbf{s}^{t+1}_{ua},I_{j})$
where $\mathcal{I}$ is the item set containing all items produced by author $a$.
3. Methodology
In this section, we propose the RLIV-UA model to represent the SCRI-MDP in the UA pair state space. Firstly, due to the long time span between adjacent states of UA pair in the SCRI-MDP, we propose the Adjacent State Approximation (ASA) method to construct RL training samples. Then, we introduce the detailed network architecture of multi-task LIV networks and the final online deployment of the RLIV-UA model. Note that the overall framework of the RLIV-UA model is shown in Fig. 3.
3.1. Adjacent State Approximation
Since adjacent UA interactions happen in requests with a long time span, and the UA state space is extremely large with sparse UA interactions, the LIV of UA pairs is modeled as the SCRI-MDP. On one hand, the next state $s^{\prime}$ in the RL training sample $(s,a,r,s^{\prime})$ is not available until the next interaction of the same UA pair happens. On the other hand, it is impossible to store the interaction traces for all UA pairs in the industrial RS. Therefore, we design that the state space only contains user or author static features and the dynamic features between them:
(3) $\mathbf{s}^{t}_{ua}=\{\mathbf{f}_{u},\mathbf{f}_{a},\mathbf{f}_{ua}^{t}\}$
Specifically, the dynamic features $\mathbf{f}_{ua}^{t}$ in the current state $\mathbf{s}^{t}_{ua}$ is defined as the counts of several kinds of interactions between $ua$. Note that $\mathbf{f}_{ua}^{t+1}$ can be approximated based on user’s current rewards $r^{t}_{ua}$ and current dynamic UA features $\mathbf{f}_{ua}^{t}$:
(4) $\mathbf{f}^{t+1}_{ua}=\begin{cases}\mathbf{f}_{ua}^{t}+1,&\text{if }r^{t}_{ua}>0\\ \mathbf{f}_{ua}^{t},&\text{otherwise }\end{cases}$
Since the SCRI-MDP only focuses on the interactions and state transitions between UA pair like $ua$, it ignores interactions between user $u$ and items of other authors. Under the above assumption of the SCRI-MDP, the next state $\hat{\mathbf{s}}^{t+1}_{ua}$ can be derived by the dynamic features $\mathbf{f}^{t+1}_{ua}$.
@@ 3.2. Multi-Task LIV Networks
In order to model the progressive changes of lifelong UA relationship, we elaborately design 4 LIV networks based on the MTCL architecture, including Click, Watch time, Follow and Gift LIV networks. In particular, denser interaction signals (e.g. click and watch time) are used to compensate for sparse labels (e.g. follow and gift). For simplicity, we take one task tower as an example to explain the network structure.
As shown in Fig. 3, the current $ua$ state $\left\{\mathbf{f}_{u},\mathbf{f}_{a},\mathbf{f}_{ua}^{t}\right\}$ is fed into a shared Embedding Layer to obtain corresponding hidden embeddings $\mathbf{h}_{u}$, $\mathbf{h}_{a}$ and $\mathbf{h}_{ua}^{t}$. Taking the vector $\mathbf{H}^{t}_{ua}=\mathrm{concat}(\mathbf{h}_{u},\mathbf{h}_{a},\mathbf{h}_{ua}^{t})$ as the network input, the $Q_{\phi}(\mathbf{H}^{t}_{ua})$ is denoted as the final LIV of $ua$ at the $t$-th ineraction. To mitigate the deviation problem of value overestimation (Van Hasselt et al., 2016), double value networks and two corresponding target networks are used to output the minimum value:
(5) $Q_{\phi}(\mathbf{H}^{t}_{ua})=min(Q_{\phi^{1}}(\mathbf{H}^{t}_{ua}),Q_{\phi^{2}}(\mathbf{H}^{t}_{ua}))$
@@
(6) $\mathcal{L}(\phi)$ $=$ $\mathbb{E}_{(\mathbf{s}^{t}_{ua},r^{t}_{ua},\mathbf{s}^{t+1}_{ua})\in D}[(Q_{\phi}(\mathbf{H}^{t}_{ua})-y)^{2}]$
(7) $y$ $=$ $r^{t}_{ua}+\gamma\max_{j\in\mathcal{I}}Q_{\phi^{\prime}}^{\prime}(\mathbf{H}^{t+1}_{ua})$
where $D$ indicates the real-time collected sample buffer, $y$ indicates the target output value of the LIV network, and $Q_{\phi^{\prime}}^{\prime}$ represents the output value of target network with same structure as LIV network $Q_{\phi}$. Note that the network parameter $\phi^{\prime}$ of $Q_{\phi^{\prime}}^{\prime}$ is periodically copied from $\phi$ of $Q_{\phi}$.
To reinforce the learning of sparse gift label of UA pairs in the Gift tower, an assistance neural network with two full-connected layers is designed to predict whether user $u$ will gift author $u$ at this interaction. As shown in Fig. 3, the user and author static hidden embeddings are input to the assistance network. And the binary cross entropy loss of the assistance gift binary classification goal is added to total loss.
@@ 3.3. Auxiliary Supervised Learning Network
Previous work (Liu et al., 2024) finds that the target value $y$ is often dominated by the inaccurate output of the target network $Q_{\phi^{\prime}}^{\prime}(\mathbf{H}^{t+1}_{ua})$ in practice, due to the instability of critic learning in RL. This problem reduces the effectiveness of the real reward $r^{t}_{ua}$ in guiding the learning of the value network, since it becomes relatively too small to provide meaningful learning signals. Furthermore, the large scale and extreme sparsity of the UA state space make the RL model even more difficult to converge.
@@
(8) $Q_{\phi}(\mathbf{H}^{t}_{ua}):=\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})+\gamma\times\hat{V}_{\phi}(\mathbf{H}^{t}_{ua})$
where $\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})$ represents the current predicted reward between $ua$ at the $t$-th interaction with $\eta$ as its trainable parameters, and $\hat{V}_{\phi}(\mathbf{H}^{t}_{ua})$ is the future cumulative reward value excluding the current reward.
As real $r^{t}_{ua}$ is available based on the $t$-th interaction between $ua$, the predicted reward $\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})$ can be learned by supervised loss. Incorporating with the aforementioned clipped double Q-learning (Fujimoto et al., 2018) shown in Eq. 5, the auxiliary supervised learning network improves the convergence of the RLIV-UA model.
Hence, the general loss of a LIV network is defined as:
(9) $\mathcal{L}_{Q}=huber\_loss(\hat{r}_{\eta}(\mathbf{H}^{t}_{ua}),r^{t}_{ua})+\sum_{2}^{k=1}(huber\_loss(Q_{\phi_{k}}(\mathbf{H}^{t}_{ua}),y)\$
where the first term denotes the loss for the auxiliary supervised learning network, and the second term denotes the original critic learning loss.
@@
where $\mathcal{L}_{A}^{g}$ is the assistance gift loss and $\mathcal{L}_{Q}^{l}$ is the loss function, defined in Eq. 9, for each label in the target label set $C$, respectively. In practice, LIV scores from different task towers can be selectively applied based on the actual optimization goal, such as improving long-term user retention or maximizing platform revenue.
@@ 3.4. Online Deployment
In real industrial RS, the online system architecture of RLIV-UA model is illustrated in Fig. 4, including the offline training process and the online serving process. In the offline training process, we first utilize the label joiner to merge the UA state features $s$, action $a$, and the corresponding immediate feedback $r$ at timestamp $t$. Then we leverage the ASA method to approximate the next state $s^{\prime}$ and obtain RL training sample $(s,a,r,s^{\prime})$. In the online serving process, the RLIV-UA model is deployed during the ranking stage of RS to influence the final ordering of candidate items, whose number $N$ is typically less than 300.
Specifically, at each request of a user, the RLIV-UA model outputs corresponding LIVs for N candidate items. The candidate items are then ranked based on their LIV scores. Finally, the top K candidate items are selected as the final item list and are exposed to the user by order. Subsequently, the user interacts with each exposure item to return the feedback signals to RS.
@@ 4. Experiments
The proposed RLIV-UA model is evaluated in an offline simulated recommendation scenario to verify its performance compared with state-of-the-art models and the effectiveness of its each part. It is also applied in two real industrial recommendation scenarios to verify its effectiveness in large-scale industrial RS through online A/B tests.

### v2

@@ 3.3. Multi-Task LIV Networks
The overall framework of the RLIV-UA model is illustrated in Fig. 4. To capture the progressive nature of lifelong user-author relationships, from initial discovery to deep engagement, we design four dedicated LIV networks based on the Multi-Task Critic Learning (MTCL) architecture: Click, Effective View, Follow, and Gift. Crucially, the positive sample rates for these signals vary drastically across tasks: approximately 3% for Click, 1% for Effective View, 0.2% for Follow, and a mere 0.01% for Gift. This extreme imbalance renders the Follow and Gift labels exceptionally sparse, making direct learning from them highly unstable and inefficient. To address this, MTCL leverages the relatively dense signals (Click and Effective View) as auxiliary tasks to bootstrap and stabilize the learning of the sparse, high-value actions (Follow and Gift), enabling the model to effectively propagate reward signals across the entire interaction hierarchy. For simplicity, we use a single task tower as an example to illustrate the network structure below.
As shown in Fig. 4, the current UA state $\left\{\mathbf{f}_{\mathrm{u}}^{t},\mathbf{f}_{\mathrm{a}}^{t},\mathbf{f}_{\mathrm{ua}}^{t},\Delta^{t}_{\mathrm{ua}}\right\}$ is fed into a shared embedding layer to obtain corresponding hidden embeddings $\mathbf{h}^{t}_{\mathrm{u}}$, $\mathbf{h}^{t}_{\mathrm{a}}$, $\mathbf{h}^{t}_{\mathrm{ua}}$ and $\delta^{t}_{\mathrm{ua}}$. Taking the vector $\mathbf{H}^{t}_{\mathrm{ua}}=\mathrm{concat}(\mathbf{h}^{t}_{\mathrm{u}},\mathbf{h}^{t}_{\mathrm{a}},\mathbf{h}_{\mathrm{ua}}^{t},\delta^{t}_{\mathrm{ua}})$ as the network input, $Q_{\phi}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})$ is denoted as the final LIV of $\mathrm{ua}$ at the $t$-th interaction. To mitigate the problem of value overestimation deviation (Van Hasselt et al., 2016), double value networks and two corresponding target networks are used to generate the minimum value:
(9) $Q_{\phi}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})=min(Q_{\phi^{1}}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}}),Q_{\phi^{2}}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}}))$
@@
(10) $\mathcal{L}(\phi)$ $=$ $\mathbb{E}_{(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}},\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}})\in D}[(Q_{\phi}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})-y)^{2}]$
(11) $y$ $=$ $r^{t}_{\mathrm{ua}}+\gamma Q_{\phi^{\prime}}^{\prime}(\mathbf{H}^{t+1}_{\mathrm{ua}},\underset{{a^{t+1}_{\mathrm{ua}}\in\mathcal{A}}}{\mathrm{argmax}}Q_{\phi}(\mathbf{H}^{t+1}_{\mathrm{ua}},a^{t+1}_{\mathrm{ua}}))$
where $D$ indicates the sample buffer collected in real time, $y$ indicates the target output value of the LIV network, and $Q_{\phi^{\prime}}^{\prime}$ represents the output value of the target network with the same structure as the LIV network $Q_{\phi}$. Note that the network parameter $\phi^{\prime}$ of $Q_{\phi^{\prime}}^{\prime}$ is periodically copied from $\phi$ of $Q_{\phi}$.
To reinforce the learning of the sparse gift label of UA pairs in the Gift tower, an assistance MLP network is designed to predict whether user U will gift author A at this interaction. As shown in Fig. 4, the user and the author static hidden embeddings $\mathbf{h}_{u},\mathbf{h}_{a}$ are input to the assistance network. The binary cross-entropy loss $\mathcal{L}_{A}^{g}$ of the assistance gift binary classification goal is added to the total loss.
(12) $\mathcal{L}_{A}^{g}=-\mathbb{E}_{\mathbf{s}^{t}_{\mathrm{ua}}\in D}\left[y_{g}\log(\hat{f}(\mathbf{h}_{u},\mathbf{h}_{a}))+(1-y_{g})\log(1-\hat{f}(\mathbf{h}_{u},\mathbf{h}_{a}))\right]$
where $y_{g}$ indicates the true value of whether user U will gift author A at this interaction and $\hat{f}(\mathbf{h}_{u},\mathbf{h}_{a})$ is the predicted value output by assistance network.
@@ 3.4. Auxiliary Supervised Learning Network
Previous work (Liu et al., 2024) finds that the target value $y$ is often dominated by the inaccurate output of the target network $Q_{\phi^{\prime}}^{\prime}$ in practice, due to the instability of critic learning in RL. This problem reduces the effectiveness of the real reward $r^{t}_{\mathrm{ua}}$ in guiding the learning of the value network, since it becomes relatively too small to provide meaningful learning signals. Furthermore, the large scale and extreme sparsity of the UA state space make the RL model even more difficult to converge.
@@
(13) $Q_{\phi}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}}):=\hat{R}_{\eta}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})+\gamma\times\hat{T}_{\xi}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})$
where $\hat{R}_{\eta}$ represents the reward prediction network with $\eta$ as its trainable parameters, $\hat{T}_{\xi}$ is the Q residual network with $\xi$ as its trainable parameters and $\phi=\{\eta,\xi\}$.
Since the real reward $r^{t}_{\mathrm{ua}}$ is available based on the $t$-th interaction between $\mathrm{ua}$, the predicted reward $\hat{R}_{\eta}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}})$ can be learned by supervised loss. Incorporating with the aforementioned clipped double Q-learning (Fujimoto et al., 2018) shown in Eq. 9, the auxiliary supervised learning network improves the convergence of the RLIV-UA model.
Hence, the general loss of an LIV network is defined as
(14) $\mathcal{L}_{Q}=MSE(\hat{R}_{\eta}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}}),r^{t}_{\mathrm{ua}})+\sum^{2}_{k=1}MSE(Q_{\phi^{k}}(\mathbf{H}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}}),y)\\$
where the first term denotes the loss for the auxiliary supervised learning network and the second term denotes the original critic learning loss which stops gradients for $\hat{R}_{\eta}$.
@@
where $\mathcal{L}_{A}^{g}$ is the assistance gift loss (defined in Eq. 12) and $\mathcal{L}_{Q}^{l}$ is the loss function (defined in Eq. 14), for each label in the target label set $C$, respectively.
@@ 3.5. Online Deployment
In industrial RS, the online system architecture of RLIV-UA model is shown in Fig. 5, including the offline training process and the online serving process. In the offline training process, we first utilize the label joiner to merge the UA state features $\mathbf{s}^{t}_{\mathrm{ua}}$, the action $a^{t}_{\mathrm{ua}}$, and the immediate feedback $r^{t}_{\mathrm{ua}}$ at timestamp $t$. Then we leverage the ASA method to approximate the next state $\mathbf{s}^{t+1}_{\mathrm{ua}}$ and obtain RL training sample $(\mathbf{s}^{t}_{\mathrm{ua}},a^{t}_{\mathrm{ua}},r^{t}_{\mathrm{ua}},\hat{\mathbf{s}}^{t+1}_{\mathrm{ua}})$. In the online serving process, the RLIV-UA model is deployed as a weight of rank score during the ranking stage of RS to influence the final ordering of candidate items, whose number $N$ is typically less than 300.
Specifically, at each request of a user, the RLIV-UA model outputs actions with highest LIV value from different task towers for candidate items. For simplicity, there only shows the application process of the action corresponding to a certain tower in Fig 5. The candidate items are then ranked by Eq. 1 based on their final scores. Finally, the top K candidate items are selected as the final item list and are exposed to the user by order. Then the user interacts with each exposure item to return feedback signals to RS.
In practice, the action $a^{t}_{\mathrm{ua}}$ with highest LIV value from different task towers can be selectively applied based on the actual optimization goal, such as improving long-term user retention or maximizing platform revenue. To explore different actions for ranking, the $\epsilon$-greedy strategy is applied online.
@@ 4. Experiments
The RLIV-UA model is evaluated in an offline simulated recommendation scenario to compare its performance with state-of-the-art models and RLIV-UA variants. It is also applied in two real industrial recommendation scenarios to verify its effectiveness in large-scale industrial RS through online A/B tests.
Table 1. Overall performance of all compared models in offline recommendation scenario.
Models Session Length Watch Time CTR Diversity
RankingModel 2.0132 59.4812 0.5948 0.0629
CQL 2.2660 58.0753 0.5125 0.1727
DQN 5.3744 125.8436 0.4074 0.5801
TD3 4.2519 79.0432 0.4258 0.4929
TD3-UA 4.9651 102.0916 0.4465 0.7161
FeedRec 6.7420 118.8468 0.4363 0.7059
RLUR 6.6810 151.0550 0.4519 0.7206
RLIV-UA(w/o MT & SL) 7.2940 188.0499 0.5229 0.7994
RLIV-UA(w/o MT) 9.2800 248.1804 0.5340 0.8746
RLIV-UA 12.8860 377.4867 0.5015 0.8827